# CityLoc: 6 DoF Localization of Text Descriptions in Large-Scale Scenes with Gaussian Representation

### v1

@@ Abstract
Localizing text descriptions in large-scale 3D scenes is inherently an ambiguous task. This nonetheless arises while describing general concepts, e.g. all traffic lights in a city. To facilitate reasoning based on such concepts, text localization in the form of distribution is required. In this paper, we generate the distribution of the camera poses conditioned upon the textual description. To facilitate such generation, we propose a diffusion-based architecture that conditionally diffuses the noisy 6DoF camera poses to their plausible locations. The conditional signals are derived from the text descriptions, using the pre-trained text encoders. The connection between text descriptions and pose distribution is established through pretrained Vision-Language-Model, i.e. CLIP. Furthermore, we demonstrate that the candidate poses for the distribution can be further refined by rendering potential poses using 3D Gaussian splatting, guiding incorrectly posed samples towards locations that better align with the textual description, through visual reasoning. We demonstrate the effectiveness of our method by comparing it with both standard retrieval methods and learning-based approaches. Our proposed method consistently outperforms these baselines across all five large-scale datasets. Our source code and dataset will be made publicly available.
Figure 1: CityLoc: Given an ambiguous text description, our method accurately localizes camera poses distribution in the large-scale city, mapping out the likely locations of descriptions like intersections, traffic lights, and urban areas.
🖂 Corresponding author: Bin Ren, bin.ren@insait.ai
@@ 1 Introduction
With the emergence of multi-modal understanding at scale, text and vision inputs are more coupled then ever. They are known to be complimentary, but yet not be available jointly at all times. Therefore, methods such as CLIP  [45], DALLE  [46], LLAVA  [34], and others  [26, 30, 59, 4, 73, 9, 43, 1, 10], learn to translate one to the other for several tasks. We are interested in performing localization in a visual environment, i.e. the visual representation of the environment is known, by merely using its text description. Our motivation for doing so comes in twofold: (i) To understand the visual locations based on the language description by humans for natural Human-robot interactions; (ii) To facilitate large-language models (LLMs) to explore the visual scenes through language tokens. We are interested in performing text-based 6 degrees-of-freedom (6DoF) localization at large scenes which are infeasible to be processed in a simple feed-forward manner within the practical means.
The text-based 6DoF localization – at the scale of the large 3D scenes and within the aforementioned motivations – is essentially an ambiguous task in most scenarios. For example, there are many scenarios of “building next to the street near a traffic light and zebra crossings". Therefore, we wish to perform localization covering all possible cases related to the given text description to be localized. More precisely, we wish to generate the distribution of camera poses conditioned upon the textual description. For this, we propose to use the diffusion-based conditional distribution learning technique, where the pose samples of the target distribution are obtained by iteratively updating the noisy poses, while conditioning upon the text features. This differentiates our method from those in  [22, 62], which focus on finding the most likely pose and require highly detailed text descriptions.
@@
For the visual representation, we use multiple large-scale 3D Gaussian scenes as shown in Fig. 2, which we built in-house using a recent hierarchical 3D Gaussian method [21]. This representation allows us to rasterize 2D images given 6DoF poses, thus allowing us to perform visual reasoning using images on the pose proposals obtained from the diffusion models. If necessary, the diffusion process itself could potentially be conditioned on the rasterized 2D images to better guide the localization process[58]. However, for the sake of efficiency –since we perform localization at the city scale – we avoid such conditioning. Instead, we utilize the visual 3D representation to further process the pose candidate obtained from the diffusion process. More precisely, the candidate poses generated by conditioning on the input text descriptions are further processed in two phases using the respective rasterized 2D images: (i) a direct filtration of the candidates and (ii) pose refinement through the optimization process, based on the similar between the input text and text description of the rasterized 2D images. Furthermore, we showcase the benefit of the proposed method for the task at hand in the presence of 2D images, as as well both image and text in the multi-modal setting.
Our method significantly differs from the closest related works: PoseDiffusion [58] and Text2Pose [22], Text2Loc [62], and RET [57]. PoseDiffusion localizes the central object, whose poses in the general settings are deterministic and less ambagious, unlike that of the text description. Furthermore, PoseDiffusion is tailored to perform multi-view bundle adjustment averaging the diffusion process and the iterative rasterization, making it not suitable the large scale localization. Note that the likelihood of finding visually similar images through rasterization decreases exponentially with the increase of the scene’s scale. On the other hand, Text2Pose, Text2Loc, and RET make the assumption that the input description is sufficient to be uniquely localized. They perform localization in 2D which hinders the possibility of visual reasoning through rasterization in our setting. In addition, they are retrieval-based and exploit the point cloud-based scene representation. In our experiments, we show that the retrieval-based solution performs poorly due to the diversity in the scene, limited textual hints, and our requirement of the 6DoF pose parameters.
@@
We apply our method across five different large-scale scenes, covering various viewpoints such as street-level and aerial views. These scenes include diverse environments, from urban to suburban areas, in both real and simulated settings, with a total area exceeding 10 square kilometers. We use Llava[34] to ensure a high diversity of text descriptions during training. During testing, we evaluate performance across different levels of text description granularity, showing that our method achieves better localization accuracy with higher granularity descriptions.
  • Experimental Setup and Benchmarking: We develop a comprehensive experimental setup designed to evaluate city-scale, text-based 6DoF localization.
  • Novel Approach for Text-Based 6DoF Localization: We propose a diffusion-based method for text-based 6DoF localization that operates effectively at the city scale.
@@
  • Pose Refinement Technique: We employ Gaussian splatting rendering for pose refinement, filtering out poorly matched poses and optimizing them by maximizing cosine similarity with text features. This guides the pose to the most relevant location for the given text description.
@@ 2 Related Work
Diffusion models [51, 14, 13], inspired by non-equilibrium thermodynamics, approximate complex data distributions by reversing a noise addition process through a series of diffusion steps. Originally developed for generative tasks, these models have demonstrated impressive results in image [14, 52, 13, 60], video [50, 15], and 3D point cloud generation [39, 70, 32, 47], as well as in natural language [3, 28] and audio generation [44]. More recently, diffusion models have been adapted for discriminative tasks, including image segmentation [2, 7] and visual grounding. However, there has been little exploration of applying diffusion models to camera localization tasks. In this work, we leverage the iterative nature of diffusion to refine probabilistic spatial representations for accurate camera pose estimation. By modeling spatial correlations through a Gaussian representation, we progressively refine the 6 DoF camera pose, achieving high precision and robustness in large-scale scenes, thus showcasing the flexibility of diffusion models in localization and scene understanding.
@@
To extend the advancements of language models [8, 55, 11, 24, 16] across modalities, Multi-Modal Large Language Models (MLLMs) integrate language and vision [66, 36, 72, 1, 71]. These models excel in tasks requiring an understanding of both text and images, which is relevant for camera localization in large-scale scenes. Flamingo [1] was among the first to align image-text pairs using gated cross-attention blocks, highlighting the importance of multi-modal integration. In 6 DoF camera localization, MLLMs enhance the interpretation of complex urban environments by integrating semantic information from both modalities. End-to-end MLLMs often fine-tune intermediate networks [23, 68] or sampler modules [67] to map visual features into the language space, improving scene representations for localization tasks. Models like BLIP-2 [27], MiniGPT-4 [72], and LLava [36] bridge modality gaps using querying transformers and two-stage training processes, demonstrating the effectiveness of combining visual and textual data to enhance localization accuracy. Other notable models such as Otter [25], mPLUG-Owl [64], and InstructBLIP [12] offer architectural inspirations. Drawing from these MLLMs, our method utilizes Gaussian representations [20] to improve 6 DoF camera localization in large-scale scenes, enhancing both accuracy and computational efficiency.
@@
Recent advancements in view synthesis, such as NeRFs [41, 5, 42, 6, 61] and 3DGS [20, 21, 37, 40, 63, 38], have revolutionized 3D scene representation by utilizing differentiable rendering and optimization. While these methods excel at modeling complex scenes, they struggle with large, unbounded 3D spaces. To address this, several methods have scaled 3D representations by dividing scenes into blocks or using multi-level decompositions, such as Mega-NeRF [56], Block-NeRF [54], CityGaussian [37], and hierarchical 3DGS [21].
Localization and mapping are closely linked, with high-quality 3D reconstructions aiding visual localization. Methods like PoseNet [18] predict camera poses but struggle in large environments. Hloc [48] improves large-scale localization by combining structure-from-motion with global features, while iNeRF [65] and iGS [53] invert NeRFs and 3DGS for 6DoF pose estimation. Text-based methods like Text2Pos [57] and Text2Loc [62] generate poses from descriptions, but can suffer from ambiguity. In contrast, our method learns a distribution of poses from text and refines them using a 3DGS map, marking the first work to bridge text descriptions with pose distributions and use Gaussian representations for large-scale pose refinement.
Figure 3: Overview of CityLoc. The left panel shows the training process, where images and multi-level of granularity text input are first converted to CLIP features. A fusion algorithm combines these features to train a pose diffusion model, mapping them to a 6DoF camera pose distribution. The right panel illustrates the inference phase, where the pose diffusion model outputs camera poses for any given text input. A pretrained Gaussian field is used to refine the poses, aligning the input text features with the rendered image features.

### v2

# CityLoc: 6DoF Pose Distributional Localization for Text Descriptions in Large-Scale Scenes with Gaussian Representation

@@ Abstract
Localizing textual descriptions within large-scale 3D scenes presents inherent ambiguities, such as identifying all traffic lights in a city. Addressing this, we introduce a method to generate distributions of camera poses conditioned on textual descriptions, facilitating robust reasoning for broadly defined concepts. Our approach employs a diffusion-based architecture to refine noisy 6DoF camera poses towards plausible locations, with conditional signals derived from pre-trained text encoders. Integration with the pretrained Vision-Language Model, CLIP, establishes a strong linkage between text descriptions and pose distributions. Enhancement of localization accuracy is achieved by rendering candidate poses using 3D Gaussian splatting, which corrects misaligned samples through visual reasoning. We validate our method’s superiority by comparing it against standard distribution estimation methods across five large-scale datasets, demonstrating consistent outperformance. Code, datasets and more information will be publicly available at our project page.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: CityLoc: Given an ambiguous text description, our method accurately estimates the camera pose distribution across a large-scale urban environment, pinpointing probable locations like parking spots. Using Vision-Language Models (VLMs), our approach also incorporates image inputs for more precise, context-aware pose localization.
With the emergence of multi-modal understanding at scale, the integration of text and visual inputs has become increasingly intertwined. Although these modalities are complementary, they are not always available jointly. Therefore, methods such as CLIP (Radford et al., 2021), DALLE (Ramesh et al., 2021), LLAVA (Liu et al., 2023c), and others (Li et al., 2024; Lin et al., 2024; Wang et al., 2023c; Bai et al., 2023; Zhu et al., 2023b; Chen et al., 2023; Peng et al., 2023), have been developed to bridge the gap between these modalities across various tasks. In this work, we aim to estimate pose distributions within a city-scale visual environment, leveraging textual descriptions, where the 3D contextual environment is already known. Our motivation for pursuing this research is twofold: (i) To decode visual locations from human language descriptions, facilitating natural human-robot interactions; (ii) To enable large language models (LLMs) to navigate and interpret visual scenes using language tokens. Our goal is to achieve text-based 6 degrees of freedom (6DoF) localization in expansive scenes where conventional feed-forward processing is impractical.
Text-based 6DoF localization in large 3D scenes, driven by our motivations, is inherently ambiguous. For instance, consider the scenario described by "building next to the street near a traffic light and zebra crossings" which can apply to numerous locations. We aim to achieve distributional localization that accounts for all potential instances that fit a given textual description, i.e., we intend to generate a distribution of camera poses conditioned on the text description.
To achieve this, we construct a dataset of pose-text pairs using Large Multimodal Model. We then propose a diffusion-based conditional distribution learning technique to iteratively refine noisy initial poses by conditioning on mixed visual-text features. Our approach stands in contrast to existing methods cited in  (Kolmet et al., 2022; Xia et al., 2024), which seek the most probable pose and depend on highly detailed textual descriptions.
@@
For visual representation, we use multiple large-scale 3D Gaussian scenes as shown in Fig. 2, built in-house using a recent hierarchical 3D Gaussian method (Kerbl et al., 2024). This allows us to rasterize 2D images from specified 6DoF poses, facilitating visual reasoning based on the outputs of diffusion models. While the diffusion process could be conditioned on these rasterized images to enhance localization precision (Wang et al., 2023b), for efficiency in city-scale tasks, we avoid direct conditioning. Instead, we refine pose candidates through a two-step approach: (i) Initial filtration of pose candidates to remove unlikely poses, and (ii) Subsequent refinement of poses by comparing the input text with descriptions from the rasterized 2D images. Our method’s effectiveness is demonstrated in handling both single-modality scenarios with only text and multi-modal settings that integrate images and text.
Our method significantly differs from existing works like PoseDiffusion (Wang et al., 2023b), Text2Pose (Kolmet et al., 2022), Text2Loc (Xia et al., 2024), and RET (Wang et al., 2023a). PoseDiffusion, designed for central objects with deterministic poses, struggles with text-driven ambiguity and scale, relying on multi-view adjustments that falter in large-scale settings. In contrast, Text2Pose, Text2Loc, and RET depend on precise textual descriptions for 2D localization, lacking visual reasoning and proving inefficient in our experiments due to diverse scenes, sparse text, and the challenges of estimating 6DoF poses.
@@
We deploy our method across 5 diverse large-scale scenes, covering over 10 square kilometers and featuring various viewpoints from street-level to aerial. These scenes include both urban and suburban environments in real and simulated formats. To ensure varied text descriptions during training, we utilize Llava (Liu et al., 2023c). During testing, we evaluate performance across varying levels of description granularity, incorporating both textual and visual inputs. Our results show that more detailed descriptions significantly enhance localization accuracy, affirming our method’s effectiveness in varied settings.
The contributions of this work are:
  1. 1.
    Experimental Setup and Benchmarking: We established a robust experimental framework to assess city-scale, text-based 6DoF pose distribution estimation, incorporating corresponding large scale Gaussian splats as 3D representation.
  2. 2.
    Novel Approach for Text-Based 6DoF Localization: We present a novel diffusion-based approach that integrates mixup training with multimodal image-text features using CLIP. This method effectively bridges the gap between textual descriptions and accurate spatial localization, achieving robust performance in large-scale urban environments.
  3. 3.
@@
    Pose Refinement Technique: We employ Gaussian splatting rendering to enhance the pose accuracy, discarding mismatches and optimizing alignment by leveraging cosine similarity with text features. This ensures the refined pose corresponds precisely to the most relevant location based on the provided text description.
  4. 4.
@@ 2 Related Work
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Dhariwal & Nichol, 2021), inspired by non-equilibrium thermodynamics, approximate complex data distributions by reversing a noise addition process through a series of diffusion steps. Originally developed for generative tasks, these models have demonstrated impressive results in image (Ho et al., 2020; Song et al., 2020; Dhariwal & Nichol, 2021; Wang et al., 2022), video (Singer et al., 2022; Ho et al., 2022), and 3D point cloud generation (Lyu et al., 2021; Zhao et al., 2024; Liu et al., 2023a; Ren et al., 2024), as well as in natural language (Austin et al., 2021; Li et al., 2022) and audio generation (Popov et al., 2021). More recently, diffusion models have been adapted for discriminative tasks, including image segmentation (Amit et al., 2021; Brempong et al., 2022) and visual grounding. However, there has been little exploration of applying diffusion models to camera localization tasks. In this work, we leverage the iterative nature of diffusion to refine probabilistic spatial representations for accurate camera pose estimation. By modeling spatial correlations through a Gaussian representation, we progressively refine the 6 DoF camera pose, achieving high precision and robustness in large-scale scenes, thus showcasing the flexibility of diffusion models in localization and scene understanding.
Figure 3: Overview of CityLoc. In the training process, where images and multi-level of granularity text input are first converted to CLIP features. A mix algorithm combines these features to train a pose diffusion model, mapping them to a 6DoF camera pose distribution. In the inference phase, where the pose diffusion model outputs camera poses for any given text input. A pretrained Gaussian representation is used to refine the poses, aligning the input text features with the rendered image features.
@@
To extend the advancements of language models (Brown et al., 2020; Touvron et al., 2023; Chowdhery et al., 2023; Le Scao et al., 2023; Hoffmann et al., 2022) across modalities, Multi-Modal Large Language Models (MLLMs) integrate language and vision (Yin et al., 2023; Liu et al., 2024b; Zhu et al., 2023a; Alayrac et al., 2022; Zheng et al., 2024). These models excel in tasks requiring an understanding of both text and images, which is relevant for camera localization in large-scale scenes. Flamingo (Alayrac et al., 2022) was among the first to align image-text pairs using gated cross-attention blocks, highlighting the importance of multi-modal integration. In 6 DoF camera localization, MLLMs enhance the interpretation of complex urban environments by integrating semantic information from both modalities. End-to-end MLLMs often fine-tune intermediate networks (Lai et al., 2024; Zhang et al., 2023) or sampler modules (You et al., 2023) to map visual features into the language space, improving scene representations for localization tasks. Models like BLIP-2 (Li et al., 2023b), MiniGPT-4 (Zhu et al., 2023a), and LLava (Liu et al., 2024b) bridge modality gaps using querying transformers and two-stage training processes, demonstrating the effectiveness of combining visual and textual data to enhance localization accuracy. Other notable models such as Otter (Li et al., 2023a), mPLUG-Owl (Ye et al., 2023), and InstructBLIP (Dai et al., 2024) offer architectural inspirations. Drawing from these MLLMs, our method utilizes Gaussian representations (Kerbl et al., 2023) to improve 6 DoF camera localization in large-scale scenes, enhancing both accuracy and computational efficiency.
@@
Recent advancements in view synthesis, such as NeRFs (Mildenhall et al., 2021; Barron et al., 2021; Müller et al., 2022; Barron et al., 2022; Wu et al., 2023) and 3DGS (Kerbl et al., 2023, 2024; Liu et al., 2024c; Ma et al., 2025; Yang et al., 2024; Lou et al., 2024), have revolutionized 3D scene representation by utilizing differentiable rendering and optimization. To address the problem of large scenes, several methods have divided scenes into blocks or using multi-level decompositions, such as Mega-NeRF (Turki et al., 2022), Block-NeRF (Tancik et al., 2022), CityGaussian (Liu et al., 2024c), and hierarchical 3DGS (Kerbl et al., 2024).
Localization and mapping are closely linked, with high-quality 3D reconstructions aiding visual localization. Methods like PoseNet (Kendall et al., 2015) predict camera poses but struggle in large environments. iNeRF (Yen-Chen et al., 2021) and iGS (Sun et al., 2023) invert NeRFs and 3DGS for 6DoF pose estimation. Text-based methods like Text2Pos (Kolmet et al., 2022) and Text2Loc (Xia et al., 2024) generate poses from descriptions, but can suffer from ambiguity. In contrast, our method learns a distribution of poses from text and refines them using a 3DGS map, marking the first work to bridge text descriptions with pose distributions and use Gaussian representations for large-scale pose refinement.
Algorithm 1 Mixup Training Algorithm
1:
2:Inputs:
  • Image embedding $\mathcal{T}_{f}$ and text embedding $\mathcal{T}_{t}$
  • Random pose $P_{\text{rand}}$ and ground truth pose $P_{\text{gt}}$
  • Swap ratio $\beta\in[0,1]$ (hyperparameter)
  • Denoiser $\mathcal{F}_{\theta}$
3:Estimate the pose distribution $p(P\mid\mathcal{T})$
4:for epoch $=1$ to $N$ do
5:     for each batch $(\mathcal{T}_{f},\mathcal{T}_{t},P_{gt})\in\mathcal{D}$ do
6:         Sample a timestamp $t\sim\text{Uniform}(\{1,\dots,T\})$
7:         Random Sample embeddings:
$\mathbf{\mathcal{T}}_{\text{mix}}\leftarrow\begin{cases}\mathcal{T}_{t}&\text{with swap ratio }\beta,\\ \mathcal{T}_{f}&\text{with swap ratio }1-\beta.\end{cases}$
8:         Sample noise and predict noise
9:         Predict pose: $P_{\text{pred}}\leftarrow h_{\text{DDPM}}(\mathbf{T}_{\text{mix}},P_{\text{rand}})$
10:         Compute loss: $\mathcal{L}\leftarrow\|P_{\text{pred}}-P_{\text{gt}}\|^{2}$
11:         Update model: $\mathcal{F}_{\theta}\leftarrow\arg\min_{h}\mathbb{E}[\mathcal{L}]$
12:     end for
13:end for
14:return $\mathcal{F}_{\theta}$